OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46 centimeters. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences.

OpenStudy (hoblos):

x+x+y = 46 y=x+4

OpenStudy (anonymous):

hoblos is on fire :D

OpenStudy (hoblos):

do you need to solve it

OpenStudy (hoblos):

lol

OpenStudy (anonymous):

An eery, blue fire.

OpenStudy (anonymous):

Yes, I forgot to put that in the question.

OpenStudy (hoblos):

x+x+x+4 = 46 3x=42 x=14 y=18

OpenStudy (anonymous):

Its' pink fsm, don't tell me that you color blind ? :P

OpenStudy (anonymous):

LOL. Thanks guys!